OpenStudy (anonymous):

Determine the domain of the following function.{(-2,3),(0,5),(3,7),(6,-3)}

OpenStudy (firejay5):

this is a math question

OpenStudy (anonymous):

This is a math question but the answer I believe is, \[D \left\{ -2 \le X \le 6\right\}\]